# ZIP: Scalable Crowd Counting via Zero‑Inflated Poisson Modeling

### v2

@@ Introduction
ZIP-B (ours) 47.81 75.04 5.51 8.63 69.46 121.88 28.26 64.84
Table 3: Comparison of our model ZIP-B with state-of-the-art approaches on ShanghaiTech A & B, UCF-QNRF, and the NWPU-Crowd validation split. ZIP-B achieves the best performance across all these four datasets under both MAE and RMSE, surpassing all strong baselines.
Effectiveness of ZIP
In this subsection, we empirically validate the effectiveness of our proposed ZIP framework by comparing it against several strong baselines, including BL (Ma et al. 2019), DMCount (Wang et al. 2020a), and EBC (Ma, Sanchez, and Guha 2024). For a fair comparison, we adopt a consistent encoder-decoder architecture based on VGG19 (Simonyan and Zisserman 2015). Before proceeding to the comparison, we first determine the optimal value of the weighting parameter $\omega$ in Eq.(11), which balances the cross-entropy term and other two terms. To this end, we evaluate the VGG19-based model under various settings of $\omega$ on the ShanghaiTech B dataset and report the results in Fig. 5. As illustrated, the model achieves the best MAE and RMSE when $\omega=0.75$, and the lowest NAE when $\omega=1.00$. In our subsequent comparisons, we fix $\omega=1.00$ to prioritize NAE, due to its robustness to varying crowd densities.
To validate the effectiveness of the proposed ZIP NLL loss, we conduct an ablation study where we replace the ZIP NLL component in the total loss function (11) with several commonly used alternatives, including MAE, DMCount (Wang et al. 2020a), and Poisson NLL. Other components in the loss function, such as the cross-entropy loss and the count loss, are kept unchanged. All models are trained using the same VGG19-based architecture under identical settings. The evaluation results on ShanghaiTech B are presented in Table 2. As shown, our ZIP NLL achieves the lowest MAE (6.36) and NAE (5.00%), and performs competitively in RMSE (10.32), demonstrating its effectiveness in enhancing counting accuracy while maintaining robust distribution modeling.
@@
Table 4: Comparison of our method ZIP-B with the latest approaches on the test split of NWPU-Crowd. ZIP achieves both the lowest MAE and NAE.
ZIP-B Compared with State-of-the-Art
We first compare our proposed model, ZIP-B, with recent state-of-the-art crowd counting methods on four benchmark datasets: ShanghaiTech A & B, UCF-QNRF, and the validation split of NWPU-Crowd. The results are summarized in Table 3. Our method consistently achieves the best performance across all datasets. In particular, on ShanghaiTech A, our model achieves the lowest MAE (47.81) and RMSE (75.04), outperforming the previous best result (48.8 MAE, 76.7 RMSE by APGCC). On ShanghaiTech B, we also achieve the best performance with a MAE of 5.51 and RMSE of 8.63, significantly surpassing strong baselines such as CrowdHat and STEERER. For the challenging UCF-QNRF dataset, our method sets a new state-of-the-art with 69.46 MAE and 121.88 RMSE. On NWPU-Crowd (val), we obtain a substantial improvement over prior methods, achieving 28.26 MAE and 64.84 RMSE, significantly outperforming the closest competitor, CLIP-EBC, which reports 36.6 MAE and 81.7 RMSE.
We further evaluate our method on the NWPU-Crowd test set, with results summarized in Table 4. Our ZIP-B achieves the lowest MAE (60.1) and lowest NAE (0.104) among all methods, indicating superior overall accuracy and robustness. Although the RMSE (299.0) is not the lowest, it is comparable to other top-performing methods. Notably, RMSE tends to amplify the influence of outlier errors. Given our observation of annotation noise, such as duplicate or out-of-bound head labels in the training and validation splits, RMSE may not always reflect true performance. In contrast, MAE and NAE offer more stable and representative evaluations, making our results more reliable in practice. These results clearly demonstrate the effectiveness and generalizability of our method across various crowd densities and scenes.
@@
MobleCount×$\times 1.25$ (Wang et al. 2020b) 5.47 34.07 82.9 137.9 8.2 13.2 124.5 207.6
ZIP-T (ours) 10.53 61.39 56.36 90.55 6.67 9.90 76.02 129.40
MobleCount×$\times 2$ (Wang et al. 2020b) 13.39 63.03 81.4 133.3 8.1 12.7 117.9 207.5
ZIP-S (ours) 33.60 242.43 55.17 88.99 5.83 9.21 73.32 125.09
Table 5: Comparison of our lightweight models with state-of-the-art lightweight approaches on ShanghaiTech A & B, and UCF-QNRF. Models are grouped by parameter size: $<1$ M, $1-10$ M, $10-20$ M, $\geq 20$ M. FLOPs are measured on $1920\times 1080$ resolution. The best result in each group is highlighted, the overall best result is shown in bold.
To investigate the performance-efficiency trade-off, we compare our proposed lightweight models (ZIP-P, ZIP-N, ZIP-T, and ZIP-S) with a range of representative compact crowd counting methods. We divide models into four groups based on parameter size: (1) ultra-lightweight models ($<$1M), (2) lightweight models (1–10M), (3) mid-size models (10–20M), and (4) larger lightweight models ($\geq$20M). Table 5 presents the details.
@@
Mid-size (10–20M). ZIP-T leads this group with the best results on all benchmarks. It achieves group-best scores of 56.36 MAE on ShanghaiTech A, 6.67 on ShanghaiTech B, and 76.02 on UCF-QNRF, significantly outperforming MobileCount×$\times$2 with fewer parameters and FLOPs.
Larger lightweight ($\geq$20M). ZIP-S sets a new benchmark in this group and across all lightweight models, achieving 55.17 MAE on ShanghaiTech A, 5.83 on ShanghaiTech B, and 73.32 on UCF-QNRF. These scores represent the best results among all compared lightweight models, and they also approach or surpass those of regular-sized models presented in Table 3.
@@
We also evaluate ZIP and DMCount on the same NWPU-Crowd validation image, using an identical VGG-19 encoder–decoder backbone for fair comparison. As shown in Fig. 6, the two density maps differ markedly. DMCount produces elongated, ribbon-like streaks inside each dense block, which is a typical artifact of its transport-based loss. In contrast, ZIP yields compact, circular kernels that more closely resemble the point annotations, preserving the grid structure of the display. As for accuracy, DMCount under-counts the scene by 114 heads (-6.5%), while ZIP misses only 19 heads (-1.1%), achieving a five-fold reduction in absolute error.
@@ Conclusion
We presented ZIP, a crowd counting framework based on zero-inflated Poisson regression. By learning a structural-zero branch to suppress background and non-head body regions, and a Poisson-rate branch to model blockwise counts, the network tackles the extreme sparsity of density maps without any explicit segmentation supervision. Comprehensive experiments show that the full model, ZIP-B, establishes new state-of-the-art MAE and RMSE on four datasets. We further introduce a family of lightweight variants (from Pico to Small) that retain these accuracy gains under tight computational budgets and consistently outperform existing models of comparable size on all four benchmarks, highlighting both the effectiveness and scalability of ZIP.
@@
For a single block with parameters $\theta\coloneqq(\pi,\lambda)$ lying in a compact set $\Theta\coloneqq[\pi_{\mathrm{min}},\pi_{\mathrm{max}}]\times[\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}]\subset(0,1)\times(0,\Lambda]$  [^1] In practice, we generate $\pi$ using sigmoid and $\lambda$ using an average of finite bin centers, and thus we have $10<\pi_{\mathrm{min}}\leq\pi_{\mathrm{max}}<$ 1 and $0<\lambda_{\mathrm{min}.}\leq\lambda_{\mathrm{max}}<\infty$, define
@@
Then, for $k\in[0,\lambda_{\mathrm{max}}-1]$, we have the $L$-steep bound:
$|\mathcal{L}_{\theta}(k+1)-\mathcal{L}_{\theta}(k)|$ $=\left|\log\frac{p_{\theta}(k)}{p_{\theta}(k+1)}\right|$
$=\left|\log\frac{k+1}{\lambda}\right|$
$\leq\sup_{k,\lambda\in[\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}]}\left|\log\frac{k+1}{\lambda}\right|$
$=\log\frac{k_{\mathrm{max}}+1}{\lambda_{\mathrm{min}}}\eqqcolon L,$
where $k_{\mathrm{max}}$ is the largest count that ever appears in one block in the training set. Note that $k_{\mathrm{max}}$ is restricted to a relatively small number to suppress label noise following EBC framework (Ma, Sanchez, and Guha 2024).
Following (Wang et al. 2020a), we flatten the spatial dimensions to avoid cluttered notations. For a full image with $n\coloneqq hw$ blocks, let
$g(\boldsymbol{z};\boldsymbol{\theta})=\frac{1}{n}\sum_{i=1}^{n}\mathcal{L}_{\boldsymbol{\theta}_{i}}(\boldsymbol{z}_{i}),$
For two annotation-prediction pairs $(\boldsymbol{z},\boldsymbol{\theta})$ and $(\boldsymbol{z}^{\prime},\boldsymbol{\theta})^{\prime}$:
$|g(\boldsymbol{z};\boldsymbol{\theta})-g(\boldsymbol{z}^{\prime};\boldsymbol{\theta}^{\prime})|$
$\leq$ $\frac{1}{n}\sum_{i=1}^{n}\left|\mathcal{L}_{\boldsymbol{\theta}_{i}}({\boldsymbol{z}_{i}})-\mathcal{L}_{\boldsymbol{\theta}_{i}}({\boldsymbol{z}_{i}^{\prime}})\right|+\frac{1}{n}\sum_{i=1}^{n}\left|\mathcal{L}_{\boldsymbol{\theta}_{i}}({\boldsymbol{z}_{i}^{\prime}})-\mathcal{L}_{\boldsymbol{\theta}_{i}^{\prime}}({\boldsymbol{z}_{i}^{\prime}})\right|$
$\leq$ $\frac{L}{n}\sum_{i=1}^{n}\left|\boldsymbol{z}_{i}-\boldsymbol{z}_{i}^{\prime}\right|+M\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\|_{1},$
@@
$M\coloneqq\sup_{k,\theta}\|\nabla_{\theta}\mathcal{L}_{\theta}(k)\|<\infty$
due to the compactness of $\Theta$.
Combining the two sums gives
$|g(\boldsymbol{z};\boldsymbol{\theta})-g(\boldsymbol{z}^{\prime};\boldsymbol{\theta}^{\prime})|$ $\leq L\|\boldsymbol{z}-\boldsymbol{z}^{\prime}\|_{1}+M\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\|_{1}$
$\leq C\|(\boldsymbol{z},\boldsymbol{\theta})-(\boldsymbol{z}^{\prime},\boldsymbol{\theta}^{\prime})\|_{1}$
where $C\coloneqq\max(L,M)<\infty$. Hence, the ZIP loss is $C$-Lipschitz under the joint $\mathcal{L}_{1}$ norm.
As in the supplement of (Wang et al. 2020a), we treat the crowd counting model as $n$ identical scalar regressors. Lemma E (b) in (Wang et al. 2020a) gives
$\mathcal{R}_{\mathcal{S}}(\mathcal{L}_{\mathrm{ZIP}}\circ\mathcal{F}^{\prime})\leq C\cdot n\cdot\mathcal{R}_{\mathcal{S}}(\mathcal{H}).$
Plugging this into Theorem 26.5 in (Shalev-Shwartz and Ben-David 2014) yields, $\forall\delta\in(0,1)$, with the probability at least $1-\delta$ over the draw of the training sample $\mathcal{S}$ of size $K$,
$\mathcal{R}_{\mathcal{D}}(f_{\mathcal{S}}^{\mathrm{ZIP}},\mathcal{L}_{\mathrm{ZIP}})\leq$ $\mathcal{R}_{\mathcal{D}}(f_{\mathcal{D}},\mathcal{L}_{\mathrm{ZIP}})+2\cdot C\cdot n\cdot R_{\mathcal{S}}(\mathcal{H})+$
@@
where $B=\sup_{\boldsymbol{X},\boldsymbol{z}}\mathcal{L}_{\mathrm{ZIP}}(\boldsymbol{z},f(\boldsymbol{X}))$. ∎
@@
Table 6: Inference speed (frames per second, FPS) of the lightweight variants on ShanghaiTech B ($1024\times 768$ resolution). Results are reported on four hardware platforms.

### v3

@@ Introduction
ZIP-B (ours) 47.8 75.0 5.5 8.6 69.4 121.8 28.2 64.8
Table 3: Comparison of our model ZIP-B with state-of-the-art models of similar sizes on ShanghaiTech A & B, UCF-QNRF, and the NWPU-Crowd validation split. ZIP-B achieves the best performance across all these four datasets under both MAE and RMSE, surpassing all strong baselines.
Ablation Study
We first determine the optimal value of the weighting parameter $\omega$ in Eq.(11), which balances the cross-entropy term and other two terms. To this end, we evaluate the VGG19-based encoder-decoder model under various settings of $\omega$ on the ShanghaiTech B dataset and report the results in Fig. 5. As illustrated, the model achieves the lowest MAE (6.03) when $\omega=1.00$, and thus, we fix $\omega=1.00$ at this value for all subsequent experiments.
To validate the effectiveness of the proposed ZIP NLL loss, we conduct an ablation study where we replace the ZIP NLL component in the total loss function in Eq. (11) with several commonly used alternatives, including MAE, MSE, DMCount (Wang et al. 2020a), and Poisson NLL. Other components in the loss function, such as the cross-entropy loss and the count loss, are kept unchanged. All models are trained using the same VGG19-based architecture under identical settings. The evaluation results on ShanghaiTech B are presented in Table 2. As shown, our ZIP NLL achieves the lowest MAE (6.03), RMSE (9.95) and NAE (4.99%), demonstrating its effectiveness in enhancing counting accuracy while maintaining robust distribution modeling.
ZIP-B Compared with State-of-the-Art
We first compare our base model, ZIP-B , with state-of-the-art crowd counting methods of comparable parameter sizes on four benchmark datasets: ShanghaiTech A & B, UCF-QNRF, and the validation split of NWPU-Crowd. The results are summarized in Table 3. Our method consistently achieves the best performance across all datasets. In particular, on ShanghaiTech A, our model achieves the lowest MAE (47.8) and RMSE (75.0), outperforming the previous best result (48.8 MAE, 76.7 RMSE by APGCC). On ShanghaiTech B, we also achieve the best performance with an MAE of 5.5 and an RMSE of 8.6, significantly surpassing strong baselines such as CrowdHat and STEERER. For the challenging UCF-QNRF dataset, our method sets a new state-of-the-art with 69.4 MAE and 121.8 RMSE. On NWPU-Crowd (val), we obtain a substantial improvement over prior methods, achieving 28.2 MAE and 64.8 RMSE, significantly outperforming the closest competitor, CLIP-EBC, which reports 36.6 MAE and 81.7 RMSE.
We further evaluate our method on the NWPU-Crowd test set, with results summarized in Table 4. ZIP-B achieves the lowest MAE (60.1) and lowest NAE (0.104). Compared to the previous best-performing models, MAE is reduced by 1.2 (from 61.3 by CLIP-EBC), and NAE is improved by 21.8% relative to STEERER (from 0.133 to 0.104). While the RMSE (299.0) is not the lowest among all methods, it remains competitive. Given that RMSE is highly sensitive to annotation noise (as illustrated in the supplementary material), MAE and NAE serve as more stable and representative metrics. These results highlight the effectiveness and robustness of ZIP-B across varying crowd densities. We also provide comparisons of ZIP-B with other models under varying illuminance and crowd sizes on NWPU-Crowd in the supplementary material.
@@
Table 4: Comparison of our method ZIP-B with the latest approaches of comparable scales on the test split of NWPU-Crowd. ZIP-B achieves both the lowest MAE and NAE, reducing NAE by 21.8% compared to the previous best (STEERER).
@@
MobleCount×1.25$\times$ 1.25× 1.25 (Wang et al. 2020b) 5.47 34.07 82.9 137.9 8.2 13.2 124.5 207.6
ZIP-T (ours) 10.53 61.39 56.3 90.5 6.6 9.9 76.0 129.4
MobleCount×2$\times$ 2× 2 (Wang et al. 2020b) 13.39 63.03 81.4 133.3 8.1 12.7 117.9 207.5
ZIP-S (ours) 33.60 242.43 55.1 88.9 5.8 9.2 73.3 125.0
Table 5: Comparison of our lightweight models with state-of-the-art lightweight approaches on ShanghaiTech A & B, and UCF-QNRF. Models are grouped by parameter size: <1<1< 1 M, 1101-101 - 10 M, 102010-2010 - 20 M, 20$\geq$ 20≥ 20 M. FLOPs are measured on 1920×10801920$\times$ 10801920 × 1080 resolution. The best result in each group is highlighted, the overall best result is shown in bold.
To investigate the performance-efficiency trade-off, we compare our proposed lightweight models (from ZIP-P to ZIP-S ) with a range of representative compact methods. We divide models into four groups based on parameter size: (1) ultra-lightweight models ($<$1M), (2) lightweight models (1–10M), (3) mid-size models (10–20M), and (4) larger lightweight models ($\geq$20M). Table 5 presents the details.
@@
Mid-size (10–20M). ZIP-T leads this group with the best results on all benchmarks. It achieves group-best scores of 56.3 MAE on ShanghaiTech A, 6.6 on ShanghaiTech B, and 76.0 on UCF-QNRF, significantly outperforming MobileCount×$\times$2 with fewer parameters and FLOPs.
Larger lightweight ($\geq$20M). ZIP-S sets a new benchmark in this group and across all lightweight models, achieving 55.1 MAE on ShanghaiTech A, 5.8 on ShanghaiTech B, and 73.3 on UCF-QNRF. These scores represent the best results among all compared lightweight models, and they also approach or surpass those of regular-sized models presented in Table 3.
@@
We also evaluate ZIP and DMCount on the same NWPU-Crowd validation image, using an identical VGG19-based backbone for fair comparison. As shown in Fig. 6, the two density maps differ remarkably. DMCount produces elongated, ribbon-like streaks inside each dense block, which is a typical artifact of its transport-based loss. In contrast, ZIP yields compact, circular kernels that more closely resemble the point annotations, preserving the grid structure of the display. As for accuracy, DMCount under-counts the scene by 114 heads (-6.5%), while ZIP misses only 19 heads (-1.1%), achieving a five-fold error reduction.
@@ Conclusion
We presented ZIP, a crowd counting framework based on zero-inflated Poisson regression. By learning a structural-zero branch to suppress background and non-head body regions, and a Poisson-rate branch to model blockwise counts, the network tackles the extreme sparsity of density maps without any explicit segmentation supervision. Comprehensive experiments show that the full model, ZIP-B, establishes new state-of-the-art performance on four datasets. We further introduce a family of lightweight variants (from Pico to Small) that retain these accuracy gains under tight computational budgets and consistently outperform existing models of comparable sizes on all four benchmarks, highlighting both the effectiveness and scalability of ZIP. Future work will extend ZIP to multimodal crowd counting.
@@
For a single spatial block with parameters $\theta\coloneqq(\pi,\lambda)$ constrained within a compact domain $\Theta\coloneqq[\pi_{\mathrm{min}},\pi_{\mathrm{max}}]\times[\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}]\subset(0,1)\times[1,\Lambda]$ , we define the per-block negative log-likelihood as follows 111 Since $\pi$ is computed by applying a sigmoid activation to the output of a neural head, it naturally lies in the open interval (0,1)(0,1)( 0 , 1 ). In practice, we can further restrict $\pi$ to a closed subinterval $[\pi_{\mathrm{min}},\pi_{\mathrm{max}}]\subset(0,1)$ by employing techniques such as output clipping, batch normalization, or weight regularization to ensure the logits remain within bounded ranges. For the Poisson rate parameter $\lambda$, it is defined as a weighted average over a finite set of positive and finite bin centers $\{b_{k}\}_{k=2}^{n}$ (refer to Eq. (4)). Consequently, we can ensure that $\lambda\in[\lambda_{\mathrm{min}},\lambda_{\mathrm{max}}]$, where $1\leq b_{2}\leq\lambda_{\mathrm{min}}\leq\lambda_{\mathrm{max}}\leq b_{n}<\infty$. :
@@
Given this formulation, we establish the following regularity property of $\mathcal{L}_{\theta}(k)$.
Lemma 1.
The per-block negative log-likelihood $\mathcal{L}_{\theta}(k)$ is $l$-Lipschitz continuous with respect to the discrete count variable $k$, for all $k\in\{0,1,\cdots,k_{\mathrm{max}}\}$ and $\theta\in\Theta$, where $\Theta\subset(0,1)\times[1,\Lambda]$ is a compact parameter space  [^2] In practice, $k_{\mathrm{max}}$ is upper-bounded by the number of pixels in each block, since it is physically implausible for more than one person to occupy a single pixel..
Proof.
Case 1: For $k=0$, we compute:
$|\mathcal{L}_{\theta}(1)-\mathcal{L}_{\theta}(0)|$
$=$ $\left|\log\left(\frac{\pi+(1-\pi)e^{-\lambda}}{1-\pi}\right)+\lambda-\log\lambda\right|$
$\leq$ $\left|\log\left(\frac{\pi+(1-\pi)e^{-\lambda}}{1-\pi}\right)\right|+|\lambda-\log\lambda|$ (14)
$=$ $l_{1}(\pi,\lambda)+l_{2}(\lambda),$ (15)
@@
$l_{1}(\pi,\lambda)\coloneqq\left|\log\left(\frac{\pi+(1-\pi)e^{-\lambda}}{1-\pi}\right)\right|$
and
$l_{2}(\lambda)\coloneqq|\lambda-\log\lambda|.$
To bound $l_{1}(\pi,\lambda)$, observe that
$0<e^{-\lambda_{\mathrm{max}}}\leq\pi+(1-\pi)e^{-\lambda}\leq 1$
and
$0<1-\pi_{\mathrm{max}}\leq 1-\pi\leq 1-\pi_{\mathrm{min}}.$
Thus, we have
$0<\frac{e^{-\lambda_{\mathrm{max}}}}{1-\pi_{\mathrm{min}}}\leq\frac{\pi+(1-\pi)e^{-\lambda}}{1-\pi}<\frac{1}{1-\pi_{\mathrm{max}}},$
which implies
$l_{1}(\pi,\lambda)\leq\max\left(\left|\log\left(\frac{e^{-\lambda_{\mathrm{max}}}}{1-\pi_{\mathrm{min}}}\right)\right|,\left|\log\left(\frac{1}{1-\pi_{\mathrm{max}}}\right)\right|\right)$
For $l_{2}(\lambda)$, note that $\lambda-\log\lambda$ is monotonically increasing for $\lambda\geq 1$, so
$1\leq\lambda_{\mathrm{\min}}-\log\lambda_{\mathrm{\min}}\leq\lambda-\log\lambda\leq\lambda_{\mathrm{max}}-\log\lambda_{\mathrm{max}},$
and therefore
$l_{2}(\lambda)\leq\lambda_{\mathrm{max}}-\log\lambda_{\mathrm{max}}.$
Combining these bounds yields:
$|\mathcal{L}_{\theta}(1)-\mathcal{L}_{\theta}(0)|\leq u_{1}+u_{2},$ (16)
where
$u_{1}\coloneqq\max\left(\left|\log\left(\frac{e^{-\lambda_{\mathrm{max}}}}{1-\pi_{\mathrm{min}}}\right)\right|,\left|\log\left(\frac{1}{1-\pi_{\mathrm{max}}}\right)\right|\right)$
and
$u_{2}\coloneqq\lambda_{\mathrm{max}}-\log\lambda_{\mathrm{max}}.$
Case 2: For $k\in[1,\lambda_{\mathrm{max}}-1]$, we have:
$|\mathcal{L}_{\theta}(k+1)-\mathcal{L}_{\theta}(k)|=\left|\log\frac{k+1}{\lambda}\right|\leq\log\left(\frac{k_{\mathrm{max}}+1}{\lambda_{\mathrm{min}}}\right),$
where $k_{\mathrm{max}}$ denotes the largest observed count in any block within the training set.
Let $u_{3}\coloneqq\log(({k_{\mathrm{max}}+1})/{\lambda_{\mathrm{min}}})$, and define the Lipschitz constant $l$ as:
$l\coloneqq\max(u_{1}+u_{2},u_{3}).$
Then, for all $k\in[0,\lambda_{\mathrm{max}}-1]$, the per-block negative log-likelihood satisfies:
$|\mathcal{L}_{\theta}(k+1)-\mathcal{L}_{\theta}(k)|\leq l,$
implying that the per-block NLL $\mathcal{L}_{\theta}$ is $l$-Lipschitz continuous.
Following Wang et al., we flatten the spatial dimensions to avoid cluttered notations. Let $s\coloneqq hw$ denote the total number of spatial blocks in a single image. Define the full-image negative log-likelihood as:
$\mathcal{L}_{\mathrm{NLL}}(\boldsymbol{z};\boldsymbol{\theta})=\frac{1}{s}\sum_{i=1}^{s}\mathcal{L}_{\boldsymbol{\theta}_{i}}(\boldsymbol{z}_{i}),$ (17)
where $\boldsymbol{z}=(\boldsymbol{z}_{1},\cdots,\boldsymbol{z}_{s})$ is the vector of discrete blockwise counts, and $\boldsymbol{\theta}=(\boldsymbol{\theta}_{1},\cdots,\boldsymbol{\theta}_{s})$ are the corresponding model parameters per block, with each $\boldsymbol{\theta}_{i}=(\pi_{i},\lambda_{i})\in\Theta$. Note that Eq. (17) is the flattened version of Eq. (7). We now state the following results.
Lemma 2.
The full-image negative log-likelihood $\mathcal{L}_{\mathrm{NLL}}(\boldsymbol{z};\boldsymbol{\theta})$ is $L$-Lipschitz continuous with respect to the joint $\mathcal{L}_{1}$ norm over $(\boldsymbol{z},\boldsymbol{\theta})$.
Proof.
Let $(\boldsymbol{z},\boldsymbol{\theta})$ and $(\boldsymbol{z}^{\prime},\boldsymbol{\theta})^{\prime}$ be two annotation-parameter pairs. Then
$\left|\mathcal{L}_{\mathrm{NLL}}(\boldsymbol{z};\boldsymbol{\theta})-\mathcal{L}_{\mathrm{NLL}}(\boldsymbol{z}^{\prime};\boldsymbol{\theta}^{\prime})\right|$
$=$ $\left|\frac{1}{s}\sum_{i=1}^{s}\left(\mathcal{L}_{\boldsymbol{\theta}_{i}}(\boldsymbol{z}_{i})-\mathcal{L}_{\boldsymbol{\theta}_{i}^{\prime}}(\boldsymbol{z}_{i}^{\prime})\right)\right|$
$\leq$ $\frac{1}{s}\sum_{i=1}^{s}\left|\mathcal{L}_{\boldsymbol{\theta}_{i}}({\boldsymbol{z}_{i}})-\mathcal{L}_{\boldsymbol{\theta}_{i}}({\boldsymbol{z}_{i}^{\prime}})\right|+\frac{1}{s}\sum_{i=1}^{s}\left|\mathcal{L}_{\boldsymbol{\theta}_{i}}({\boldsymbol{z}_{i}^{\prime}})-\mathcal{L}_{\boldsymbol{\theta}_{i}^{\prime}}({\boldsymbol{z}_{i}^{\prime}})\right|$
By Lemma 2, the per-block loss $\mathcal{L}_{\theta}(k)$ is $l$-Lipschitz in $k$, so the first term is bounded by:
$\frac{1}{s}\sum_{i=1}^{s}l|\boldsymbol{z}_{i}-\boldsymbol{z}_{i}^{\prime}|=l\cdot\frac{1}{s}\|\boldsymbol{z}-\boldsymbol{z}^{\prime}\|_{1}.$
For the second term, since $\Theta$ is compact and $\mathcal{L}_{\theta}(k)$ is differentiable in $\theta$, we define
$M\coloneqq\sup_{k,\theta}\|\nabla_{\theta}\mathcal{L}_{\theta}(k)\|<\infty,$
which yields
$|\mathcal{L}_{\boldsymbol{\theta}_{i}}(\boldsymbol{z}_{i}^{\prime})-\mathcal{L}_{\boldsymbol{\theta}_{i}^{\prime}}(\boldsymbol{z}_{i}^{\prime})|\leq M\cdot\|\boldsymbol{\theta}_{i}-\boldsymbol{\theta}_{i}^{\prime}\|_{1}.$
Summing over all $i$, we obtain
$\frac{1}{s}\sum_{i=1}^{s}\left|\mathcal{L}_{\boldsymbol{\theta}_{i}}({\boldsymbol{z}_{i}^{\prime}})-\mathcal{L}_{\boldsymbol{\theta}_{i}^{\prime}}({\boldsymbol{z}_{i}^{\prime}})\right|\leq M\cdot\frac{1}{s}\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\|_{1}.$
Combining both terms gives:
$\left|\mathcal{L}_{\mathrm{NLL}}(\boldsymbol{z};\boldsymbol{\theta})-\mathcal{L}_{\mathrm{NLL}}(\boldsymbol{z}^{\prime};\boldsymbol{\theta}^{\prime})\right|\leq\frac{l}{s}\|\boldsymbol{z}-\boldsymbol{z}^{\prime}\|_{1}+\frac{M}{s}\|\boldsymbol{\theta}-\boldsymbol{\theta}^{\prime}\|_{1}$
Thus,
$|g(\boldsymbol{z};\boldsymbol{\theta})-g(\boldsymbol{z}^{\prime};\boldsymbol{\theta}^{\prime})|\leq L\|(\boldsymbol{z},\boldsymbol{\theta})-(\boldsymbol{z}^{\prime},\boldsymbol{\theta}^{\prime})\|_{1},$
where $L\coloneqq(l+M)/s<\infty$. ∎
As in the supplement of (Wang et al. 2020a), we treat the crowd counting model as consisting of $s$ independent scalar regressors, one for each spatial block in the image. Let $\mathcal{S}\coloneqq\{(\boldsymbol{X}^{(k)},\boldsymbol{z}^{(k)})\}_{k=1}^{K}$ denote a random draw of $K$ training examples from the training set $\mathcal{D}$. Let $\mathcal{F}^{\prime}$ denote the hypothesis class that mapping an image $\boldsymbol{X}$ to a vector of blockwise distribution parameters $\boldsymbol{\theta}$. Let $\mathcal{L}_{\mathrm{NLL}}$ be the ZIP loss aggregation function defined in Eq. (17). By Lemma E(b) in (Wang et al. 2020a), if each $\mathcal{L}_{\mathrm{NLL}}$ is $L$-Lipschitz continuous with respect to its input, then the empirical Rademacher complexity of the composed class is bounded as
$\mathcal{R}_{\mathcal{S}}(\mathcal{L}_{\mathrm{NLL}}\circ\mathcal{F}^{\prime})\leq L\cdot s\cdot\mathcal{R}_{\mathcal{S}}(\mathcal{H}),$
where $\mathcal{H}$ is the shared hypothesis class for the base feature extractor and prediction head across all blocks, and $\mathcal{R}_{\mathcal{S}}$ denotes the empirical Rademacher complexity with respect to sample $\mathcal{S}$. Now, applying Theorem 26.5 from (Shalev-Shwartz and Ben-David 2014) to the composed function $\mathcal{L}_{\mathrm{NLL}}\circ f^{\mathrm{ZIP}}_{\mathcal{S}}$, we obtain the following generalization bound. For any $\delta\in(0,1)$, with probability at least $1-\delta$ over $\mathcal{S}$, the expected risk is bounded by:
$\mathcal{R}_{\mathcal{D}}(f_{\mathcal{S}}^{\mathrm{NLL}},\mathcal{L}_{\mathrm{NLL}})\leq$ $\mathcal{R}_{\mathcal{D}}(f_{\mathcal{D}}^{\mathrm{NLL}},\mathcal{L}_{\mathrm{NLL}})+2\cdot L\cdot s\cdot R_{\mathcal{S}}(\mathcal{H})+$
$5B_{\mathrm{NLL}}\sqrt{\frac{2\log(8/\delta)}{K}},$
where
  • $f^{\mathrm{NLL}}_{\mathcal{S}}$ is the model learned from training sample $\mathcal{S}$,
  • $f_{\mathcal{D}}^{\mathrm{NLL}}$ is the population risk minimizer under the loss $\mathcal{L}_{\mathrm{NLL}}$,
  • $L$ is the Lipschitz constant of the loss function $\mathcal{L}_{\mathrm{NLL}}$,
  • $s$ is the number of blocks per image,
  • $B_{\mathrm{NLL}}=\sup_{\boldsymbol{X},\boldsymbol{z}}\mathcal{L}_{\mathrm{NLL}}(\boldsymbol{z},f(\boldsymbol{X}))<\infty$ is the uniform upper bound on the ZIP NLL loss, guaranteed due to the compactness of $\Theta$ and the bounded support of the count vector $\boldsymbol{z}$.
Remark.
Since the cross-entropy loss in Eq. (5) is 1-Lipschitz under the $\mathcal{L}_{1}$ norm, its empirical Rademacher complexity is bounded by $s_{+}\cdot\mathcal{R}_{\mathcal{S}}(\mathcal{H})$, where $s_{+}$ denotes the number of blocks with nonzero counts (i.e., $\boldsymbol{z}_{i}>0$) and $0\leq s_{+}\leq s$.
Combining this with the generalization bound for the counting loss in Wang et al. (2020a), we obtain the following generalization error bound for the total loss in Eq. (11):
$\mathcal{R}_{\mathcal{D}}(f_{\mathcal{S}}^{\mathrm{total}},\mathcal{L}_{\mathrm{total}})\leq$ $\mathcal{R}_{\mathcal{D}}(f_{\mathcal{D}}^{\mathrm{total}},\mathcal{L}_{\mathrm{total}})+$
$2\cdot(L\cdot s+s+\omega s_{+})\cdot R_{\mathcal{S}}(\mathcal{H})+$
@@
where $B=B_{\mathrm{NLL}}+B_{\mathrm{CE}}+B_{\mathrm{count}}$ denotes the sum of the uniform upper bounds of Eq. (7), Eq. (5) and Eq. (10).
@@
Table 6: Inference speed (frames per second, FPS) of the lightweight variants on ShanghaiTech B (1024×7681024$\times$ 7681024 × 768 resolution). Results are reported on four hardware platforms.
Dataset Training Size Block Size Scale Range Brightness Contrast Saturation Saltiness Spiciness
ShanghaiTech A 224 16 [0.75, 2.0] 0.15 0.15 0.10 0.001 0.001
ShanghaiTech B 448 16 [0.75, 2.5] 0.15 0.15 0.10 0.001 0.001
UCF-QNRF 672 32 [0.75, 2.0] 0.15 0.15 0.10 0.001 0.001
NWPU-Crowd 672 16 [0.75, 2.0] 0.20 0.20 0.15 0.001 0.001
Table 7: Dataset-specific data augmentation settings used during training. Each dataset is augmented using a tailored configuration of training resolution, block granularity, and perturbation strengths. The cropping scale controls spatial zoom variation, while brightness, contrast, and saturation simulate photometric noise. Saltiness and spiciness introduce localized binary corruption to enhance robustness.